Astronomy Report Dec 25th

Astronomy Report

Name: Federico Arturo Martinez Espinoza

Crew: Crew 201 MEx-1

Date: Dec 25 2018

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested MDRS-WF

Objects Viewed: Andromeda Galaxy

Problems Encountered: None, I used the image in Photoshop to give it color, used RGB filter to take 3 images with an exposure time of 60 s.

Operations Report Dec 25th

Crew 201 Operations Report 25-Dec-2018

SOL: 10

Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None

Generator (hours run): 17:05 yesterday until around 8: 06. Today started at 18:26

Solar— SOC 79%

Diesel Reading – Around 75% (We saw that we had a wrong measurement the last 2 days).

Propane Reading – 72%

Ethanol Free Gasoline – N/A gallons.

Water (Axillary tank) – 0 gallons

Water (static tank) – Medium level, around 185-gallons

Axillary to Static tank transfer– no

Gallons transferred: 0

Water in GreenHab – Medium level (43%)

Water (loft) – Static to Loft Pump used – yes

Water Meter: 01394934

Toilet tank emptied: NO

Deimos rover used: no, still not functional

Hours: 0

Beginning charge: –

Ending charge: –

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging X :

Spirit rover used: NO

Hours: 63

Beginning charge: 100%

Ending charge: 100%

Currently charging: YES

Opportunity rover used: NO

Hours: 44.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: YES

Curiosity rover used: NO

Hours: 61.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: YES

Notes on rovers: Yesterday we sent a bad information with the Curiosity info. It was 71% at the end of the EVA.

# Hours the ATVs were Used today: Nothing to report.

Notes on ATVs : Nothing to report.

HabCar used and why, where? Not used.

CrewCar used and why, where? Not used.

General notes and comments: We realized that we have 2 water leak in the kitchen comming from the extractor. This because of the rain .

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: – Nothing to report.

Summary of ScienceDome operations: – Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Musk Solar observatory – Nominal. RObotic Observatory – Nominal.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

Thank you!!!

Sol Summary Report Dec 25th

Crew 201 Sol Summary Report 25-Dec-2018

Sol: 10

Summary Title: A Red (and a little white) Christmas

Author’s name: Genaro Grajeda, HSO

Mission Status: Projects moving along as expected. A little Christmas cooking for a nice meal. Santa Claus showed up at the hab

Sol Activity Summary: Different methods of 3D printing, machines and filaments. Methodologies to do lasagna.

Look Ahead Plan: More 3d printing and programing, more planting, video making for educational purposes back on Earth.

Anomalies in work: Water leak due to rain informed in operations report.

Weather: Cold, some snow showers and rain.

Crew Physical Status: A little tired but everyone ok.

EVA: No EVA performed today.

Reports to be filed: Operations, Green Hab Astronomy, Photos.

Support Requested: None